OpenStudy (anonymous):

How much money will I have if I invest $1000 at 5% annual interest for ten years given the compounding?Round the answer to the nearest cent. find anually, continusly and monthly

OpenStudy (rulnick):

$1000(1.05)^10 = $1628.89

OpenStudy (anonymous):

i got 16,43.62 for anually and 1648.72 contiunusly and 10509.45 for montly but dont know if they are right

OpenStudy (rulnick):

answer above is for initial question (annual compounding)

OpenStudy (anonymous):

so are my answeres wrong

OpenStudy (rulnick):

I think the first answer (for annual) is a little high. Did you use 1000(1.05)^10?

OpenStudy (rulnick):

For monthly it would be $1000(1+0.05/12)^120 = $1647.01.

OpenStudy (rulnick):

OK?

OpenStudy (anonymous):

wouldnt it be continusly

OpenStudy (rulnick):

No, continuously I haven't done yet. Two down, one to go. Do you understand the first two?

OpenStudy (anonymous):

yes

OpenStudy (rulnick):

The continuous one is $1000e^0.5 = $1648.72.